• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Olivier

    Log entry time 10:24:06 on January22,2004

    Entry number 116414

    keyword=OTR and controlled access

    After the multiwire harp scan, MCC checked the energy spread using the OTR (see ELOG 1191570).

    We are now in controlled access (started at 9:20) to cool down the left septum.

    Bodo will check A2 in the trigger, as the rate in T6 and T7 are abnormally small compared to T5 (T7=100, T5=4500 on H2O).